Hart Enterprises recently paid a dividend, D0, of $3.25. It expects to have nonconstant growth of 20% for 2 years followed by a constant rate of 5% thereafter. The firm's required return is 19%. What is the horizon or terminal value?
Answer #1

Year 1 dividend = 3.25 (1 + 20%) = 3.9

Year 2 dividend = 3.9 (1 + 20%) = 4.68

Year 3 dividend = 4.68 (1 + 5%) = 4.914

Horizon value = D3 / required rate - growth rate

Horizon value = 4.914 / 0.19 - 0.05

Horizon value = 4.914 / 0.14

Horizon value = $35.10